write a linear equation of a line that passes through (-2,1) and is perpendicular to y= 4x-11
Answer by Alan3354(69443)

Do it like this one:
----
write a linear equation of a line that passes through (8,-6) and is perpendicular to 2x-y= 4
------------
Find the slope of the line.
Put the eqn in slope-intercept form y = mx + b. That means solve for y.
2x-y= 4
y = 2x - 4 --> m = slope = 2
---
The slope of lines perpendicular is the negative inverse = -1/2
---
y-y1 = m*(x-x1)
y + 6 = (-1/2)*(x - 8)
2y + 12 = -x + 8
x + 2y = -4